### v2

@@ 4.3 Scale and Point Systems
This study further investigates the relationship and trend between perceptual stances and resource diversity regarding GenAI across the top 100 universities. For this purpose, we developed a scale system in order to quantify different universities’ perceptions, from proactive embrace to cautious hesitance (see Table 3). The perception points are assigned according to each university’s policy stances, with ’0’ representing ’Undecided/Unclear’ policies to reflect a neutral and open position towards the use of GenAI. Various negative and positive scores in the range of [-5, 5] illustrate the spectrum from cautious hesitance to strong endorsement respectively.
@@
University policies and decisions Perception scale points
@@
[…]
Another scoring system extends to quantify the comprehensiveness of resources provided by these institutions. This is achieved by evaluating the breadth and depth of the resources, considering the target audience, the variety of resource types, and the range of content categories provided in Table 2. Specifically, universities accumulate scores based on their available resources across the three dimensions. For instance, for the target audience, a university earns one point for each distinct group for whom resources are provided, including students, faculty, or the general audience, with a maximum of three points available in this dimension. For the types of resources, a university gains one point for each type they provide, such as syllabus templates and workshops, allowing for up to five points in this dimension. Similarly, regarding the content categories, universities are awarded one point for each covered topic or theme, such as technical introductions and ethical considerations, for up to seven points for this dimension. Both quantitative scale and point systems are used specifically in Section 5.3 for exploring the relationship and trend between perceptions and resources.
@@ 5 Results
5.1 University Policies
5.1.1 Perceptions Based on Policies and Statements
Fig. 2 summarizes policies and stances adopted by the top 100 universities in the US regarding the legality and application of ChatGPT and other GenAI tools in higher educational contexts. None (0%) of the top 100 universities has completely banned these tools, reflecting a general acceptance or openness towards GenAI. The majority (54.8%) give this decision-making agency to individual instructors, indicating a contextualized and faculty-centric approach. Meanwhile, a modest 9.6% had implement conditional use policies with proper citations, and 35.6% remain either undecided or had not clearly announce their policies or stance. These responses demonstrate a diverse but flexible approach to integrating AI in higher education contexts in general.
However, it should be highlighted that no decision or no clear policy does not imply indifference toward GenAI on the part of these universities. Instead, many of them often present an open and objective introduction of ChatGPT and/or other AI writing tools, which represent their neutral perceptions. For example, the University of Illinois at Urbana-Champaign outline both the benefits and challenges of using GenAI and advise that careful thoughts and considerations should be kept in mind to incorporate GenAI into coursework. While the universities are not refusing AI tools completely, they often stand in a neutral position and share all resources in a balanced way.
When selecting the ”instructor decides” policy towards GenAI, a more cautious trend emerges, as illustrated in Fig 3 Among the 57 “instructor decides” universities, 27 adopt a stance of Prohibition by Default, only allowing the use of such tools when an instructor explicitly permits it. If the instructor has not presented any policy statements on the use of GenAI, using ChatGPT in homework and essays is generally not allowed and may be under the circumstance of plagiarism. If the instructor allows it, students must cite appropriately and take responsibility for their responses. This option reveals the universities’ more cautious perceptions and evident concerns about the impact of GenAI on academic integrity.
@@ 5.1.1 Perceptions Based on Policies and Statements
29 universities adopt a more open and neutral stance, granting instructors the autonomy to decide and addressing transparency in policy-making. This approach signals a more balanced and pragmatic perspective towards GenAI tools and also reflects the universities’ practical considerations and respect for the diverse needs and contexts of different disciplines. University of California, Irvine (UCI)’s statement serves as an example, showing the essence and rationale behind this approach.
@@
5.1.2 Purposes and Focuses of the Policies
Fig. 5 introduces the purposes and focuses of the existing policies on the use of ChatGPT in higher education. The data reveals a focus on addressing educational challenges and concerns, with higher attention to issues such as plagiarism (38%), inadequate proper attribution and citations (44%), and the limitations of AI tools (40%). On the other hand, topics related to professional research writing, such as intellectual property (20%) and data privacy (32%), have received comparatively less attention in these policies. This trend reveals that policy development in higher education institutions across the U.S. may often pay more attention to educational areas. Professional research writing and publication in academia may need more guidelines from institutions.
Figure 5: Purposes and focuses of the policies.
5.2 Guidelines and Resources
@@ 5.2.1 Target Audience
In regards to the guidelines and resources about ChatGPT and other GenAI, we first analyzed their aimed target audience (see Fig. 6). The results show that 84 universities of the top 100 universities across the US have resources and guidelines explicitly designed for faculty and instructors. A smaller portion of 23 universities offer resources aimed at students, and 19 provide guidance for the broader audience, including faculty, students, and staff, without specifying a particular audience. The findings show a broader emphasis on resources crafted for the population of faculty and instructors to incorporate AI tools into their teaching practice. Relatively fewer resources address and guide students on the appropriate application of AI tools in their learning. This discrepancy highlights the demands to further develop in-depth guidelines for students and establish a more inclusive AI resource and support system.
Figure 6: Policies from the top 100 US universities regarding GenAI.
@@ 5.2.2 Types of Resources
Fig. 7 illustrates the diversity of resources from the top 100 universities, highlighting Shared Articles or Blogs being the most prevalent type. The resource centers often feature additional reading at the bottom of the pages and/or embedded links within the content. The resources include research papers, news articles, other university websites, and blogs and covered a wide range of topics, such as opportunities and challenges of AI for education, educators’ and students’ reactions to AI innovations, and teaching strategies with the use of ChatGPT. These resources are important for instructors to gain foundational knowledge about emerging technologies.
@@
It appears from our data that 65 universities have offered syllabus templates and examples as references helping instructors and teacher trainers make policy decisions according to their own teaching contexts. These templates typically showcase three distinct policy perspectives—restrictive, mixed, and encouraging, to accommodate diverse disciplines’ contexts. Harvard University’s syllabus samples [^3] https://oue.fas.harvard.edu/ai-guidance exemplify this approach (see Fig. 8). Universities usually encourage all instructors to explicitly include a clear policy in course syllabi regarding the use and misuse of ChatGPT and other AI tools. Open and explicit communications are crucial to help students understand the boundaries and expectations when they interact with GenAI tools in learning.
@@
Additionally, 41 universities include one-on-one consultations as a type of resource for instructors and/or students. They are conducted with the institution’s AI specialists and/or representatives from the Teaching and Learning Center to address attendees’ specific concerns and navigate the personal applications of ChatGPT and other AI tools in their own teaching and learning contexts.
Furthermore, it is evident in Fig. 7 that a smaller proportion of universities offer workshops and discussions regarding the use of ChatGPT and other AI tools. This might be because many of these workshops and discussions are a part of internal resources and are not publicly available. As ChatGPT is a new emerging technology, workshops and discussions are crucial for familiarizing educators with its positive applications and educational implications [7]. The results could reflect that there is a possible demand for increasing the frequency and accessibility of these events.
@@ 5.2.3 Content Analysis of Resources and Guidelines
Fig. 9 summarizes the focuses and purposes of the existing resources and guidelines regarding the use of ChatGPT and other GenAI tools in higher education. The majority of the universities start by introducing some general technical information, such as a beginning section named What is ChatGPT? or Generative AI in Education. This approach indicates the institutions’ intention to familiarize faculty, students, and staff with AI tools and enhance their understanding of these technologies. 62 universities discuss the ethical implications of implementing ChatGPT in higher education, including their apprehensions about the misuse of AI to foster plagiarism, violate academic integrity, and negatively affect student evaluation. This discussion highlights the necessity for instructors to monitor and guide the use of ChatGPT [13], as well as the need to adapt teaching methodologies to prevent possible cheating and plagiarism.
@@
56 universities explicitly list the inherent limitations of the current Artificial Generative Intelligence (AGI) model and 38 universities address the data privacy issue, which serve as a reminder for individuals to keep in mind that these constraints may influence teaching approaches, learning experiences, and the conduction of research. Being aware of these limitations and issues ensures that such an AGI model would be used in a more responsible and ethical manner in education [17]. The limitations primarily include:
@@
The inevitable intertwining with GenAI tools needs teachers and educators to evolve the teaching methods across all subjects. Therefore, numerous top 100 universities have offered resources on pedagogical applications and prevention strategies. They cover the effective incorporation of ChatGPT into classrooms to enhance student learning experiences, alongside strategies to avoid its misuse by students. Table 4 provides a summary of major pedagogical applications and prevention strategies sourced from these leading universities. Both approaches highlight the significance of cultivating students’ critical thinking skills and problem-solving abilities, which are essential competencies in today’s rapidly advancing, technology-infused world.
@@
Pedagogical Applications Effectively incorporate AI in teaching and learning
Prevention Strategies Prevent inappropriate use of AI in teaching and learning
  • Ask students to analyze and evaluate AI-generated texts.
  • Ask students to compare and evaluate the different versions of texts generated by different AI tools.
  • Ask students to compare/contrast AI-generated texts with human writing.
  • Ask students to revise and edit AI-generated information.
  • Ask students to debate or argue with AI and reflect on their learning.
  • Use AI as a resource for students to receive feedback on their drafts
  • Use AI tools to brainstorm initial teaching ideas and activities.
  • Use AI tools to generate additional examples of certain concepts.
  • Use AI tools to summarize long or difficult text.
  • Use AI tools to generate writing prompts, grading rubrics or quiz questions based on the course materials.
  • Ask students to explain their thought processes as they solve problems.
  • Ask students to reflect on their personal learning experiences and opinions.
  • Ask students to connect with their personal knowledge and life experiences.
  • Ask students to include and provide proper academic citations.
  • Ask students to reference class materials, notes, or sources that are unavailable online.
  • Ask students to complete assignments in class.
  • Ask students to present their answers in multimodal ways, such as hand drawing, or audio threads.
  • Include visual prompts in assignments.
  • Design assignments related to current events or discussions in the specific academic field.
  • Divide the larger project into multiple smaller tasks.
@@
Another trend that emerges from the data is the discussion of using AI detection tools or AI detectors to identify AI-generated text in students’ work. 59 universities discuss the available common detective tools, such as Turnitin and GPTZero. However, it is worth noting that none (0%) of the universities in this study view the use of detective tools as a completely reliable method to identify AI-generated information, and none (0%) of them support instructors to use of the tools to evaluate students’ academic integrity and determine plagiarism. While AI detectors are designed to identify AI-generated language patterns, the research conducted by [33] show that they are not reliable in many real-world scenarios. This is particularly evident when the detective tools are faced with paraphrasing attacks which refer to applying a light paraphraser to generated texts. Even a minor rephrasing can significantly affect the performance and accuracy of the entire detection system [33]. Additionally, universities in this study raise further concerns regarding the use of AI detectors, including the potential violation of students’ intellectual property rights and the risk to data privacy once their work is submitted to the detective tools. Some universities also believe that relying on such tools might undermine the relationship of trust between students and teachers as well.
@@
5.3 Relationship Between Perceptions and Resources
5.3.1 Perception and Available Resources Across Different Tiers
This section employs the scale and point system described in method Section 4.3 to examine the relationship between three dimensions, including perceptual stances, resource diversity, and university rankings. We divide the top 100 universities into three tiers to examine the disparities in perceptions and resources across different tiers of higher education institutions. Tier 1 includes the top 1-33 universities; Tier 2 comprises the subsequent top universities ranked 34-66; and Tier 3 includes the remaining reputational universities ranked 67-100.
In terms of the results, Fig. 10 illustrates the correlation between the perceptions and resources regarding GenAI across the three ranking tiers. We find that a majority of the top 100 universities have developed a diversity of resources and guidelines for integrating ChatGPT, demonstrating a proactive approach towards applying this technology to their education. However, the figure shows that there is no significant correlation between the universities’ rankings and the depth of perceptions and resources related to ChatGPT. The trend also indicates that a large number of universities present unclear and cautious perceptions toward ChatGPT. These perceptions might be due to uncertainties and controversial features of AI in higher education.

### v3

@@ 4.3 Scale and Point Systems
This study further delves into how school ranking tiers and academic specializations shape the trends in perceptions and resource provision for GenAI in higher education. For this purpose, we developed a scale system in order to quantify different universities’ perceptions, from proactive embrace to cautious hesitance (see Table 3). The perception points are assigned according to each university’s policy stances, with ’0’ representing ’Undecided/Unclear’ policies to reflect a neutral and open position towards the use of GenAI. Various negative and positive scores in the range of [-5, 5] illustrate the spectrum from cautious hesitance to strong endorsement respectively.
@@
University policies and decisions Scale points
@@
Another scoring system extends to quantify the comprehensiveness of resources provided by these institutions. This is achieved by evaluating the breadth and depth of the resources, considering the target audience, the variety of resource types, and the range of content categories provided in Table 2. Specifically, universities accumulate scores based on their available resources across the three dimensions. For instance, for the target audience, a university earns one point for each distinct group for whom resources are provided, including students, faculty, or the general audience, with a maximum of three points available in this dimension. For the types of resources, a university gains one point for each type they provide, such as syllabus templates and workshops, allowing for up to five points in this dimension. Similarly, regarding the content categories, universities are awarded one point for each covered topic or theme, such as technical introductions and ethical considerations, for up to seven points for this dimension. Both the quantitative scale and point systems are used specifically in Section 5.3 for exploring how key variables (ranking and academic specialization) affect the trends in perceptions and resource provision.
@@ 5 Results
5.1 Policies, Management Approaches, and Perceptions from U.S. Universities
To answer RQ1, this section introduces the current policies, and management approaches adopted by the top-ranked U.S. universities, as well as the perceptions and implications derived from these policy statements. Fig. 2 summarizes policies and stances adopted by the top 100 universities (n = 104) in the U.S. regarding the legality and application of ChatGPT and other GenAI tools in higher educational contexts. None (n = 0, 0%) of the top 100 universities have completely banned these tools, reflecting a general acceptance or openness towards GenAI. The majority (n = 57, 54.8%) give this decision-making agency to individual instructors, indicating a contextualized and faculty-centric approach. Meanwhile, a modest 9.6% (n = 10) had implemented conditional use policies with proper citations, and 35.6% (n = 37) remained either undecided or had not clearly announced their policies or stance. These responses demonstrate a diverse but flexible approach to integrating AI in higher education contexts in general.
However, it should be highlighted that no decision or no clear policy does not imply indifference toward GenAI on the part of these universities. Instead, many of them often present an open and objective introduction of ChatGPT and/or other AI writing tools, which represent their neutral perceptions. For example, the University of Illinois at Urbana-Champaign outlines both the benefits and challenges of using GenAI and advises that careful thoughts and considerations should be kept in mind to incorporate GenAI into coursework. While the universities are not refusing AI tools completely, they often stand in a neutral position and share all resources in a balanced way.
When selecting the "instructor decides" policy towards GenAI, a more cautious trend emerges, as illustrated in Fig 3. Among the 57 “instructor decides” universities, 27 (47.4%) adopt a stance of Prohibition by Default, only allowing the use of such tools when an instructor explicitly permits it. If the instructor has not presented any policy statements on the use of GenAI, using ChatGPT in homework and essays is generally not allowed and may be under the circumstance of plagiarism. If the instructor allows it, students must cite appropriately and take responsibility for their responses. This option reveals the universities’ more cautious perceptions and evident concerns about the impact of GenAI on academic integrity.
@@
29 (50.9%) universities adopt a more open and neutral stance, granting instructors the autonomy to decide and addressing transparency in policy-making. This approach signals a more balanced and pragmatic perspective towards GenAI tools and also reflects the universities’ practical considerations and respect for the diverse needs and contexts of different disciplines. University of California, Irvine (UCI)’s statement serves as an example, showing the essence and rationale behind this approach.
@@
Figure 4: Policies from the top 100 US universities regarding GenAI.
@@
We are also interested in exploring the main topics covered in the existing policies concerning GenAI technologies. Fig. 5 introduces the purposes and focuses of the existing policies on the use of ChatGPT in higher education. The data reveals a focus on addressing educational challenges and concerns, with higher attention to issues such as plagiarism (n = 35, 33.7%), inadequate proper attribution and citations (n = 38, 36.5%), and the limitations of AI tools (n = 31, 29.8%). On the other hand, topics related to professional research writing, such as intellectual property (n = 13, 12.5%) and data privacy (n = 28, 26.9%), have received comparatively less attention in these policies. This trend reveals that policy development in higher education institutions across the U.S. may often pay more attention to educational areas. Professional research writing and publication in academia may need more guidelines from institutions.
Figure 5: Purposes and focuses of the policies.
@@
5.2 Guidelines and Resources for the Applications of GenAI
RQ2 asks about resources and guidance that U.S. universities provide regarding the use of GenAI tools in higher education contexts. This section reports the target audience, types of resources, and content analysis with a focus on limitations, pedagogical applications, and prevention strategies.
@@ 5.2.1 Target Audience
In regards to the guidelines and resources about ChatGPT and other GenAI, we first analyzed their aimed target audience (see Fig. 6). The results show that 70 (67.4%) universities of the top 100 universities across the U.S. have resources and guidelines explicitly designed for faculty and instructors. A smaller portion of 19 (17.8%) universities offer resources aimed at students, and 15 (14.7%) provide guidance for the broader audience, including faculty, students, and staff, without specifying a particular audience. The findings show a broader emphasis on resources crafted for the population of faculty and instructors to incorporate GenAI tools into their teaching practice. Relatively fewer resources address and guide students on the appropriate application of GenAI tools in their learning. This discrepancy highlights the demands to further develop in-depth guidelines for students and establish a more inclusive GenAI resource and support system.
@@ 5.2.2 Types of Resources
Fig. 7 illustrates the diversity of resources from the top 100 universities (n = 104), highlighting Shared Articles or Blogs being the most prevalent type (n= 74, 71.1%). The resource centers often feature additional reading at the bottom of the pages and/or embedded links within the content. The resources include research papers, news articles, other university websites, and blogs and cover a wide range of topics, such as opportunities and challenges of GenAI for education, educators’ and students’ reactions to GenAI innovations, and teaching strategies with the use of ChatGPT. These resources are important for instructors to gain foundational knowledge about emerging technologies.
@@
It appears from our data that 65 (62.5%) universities have offered syllabus templates and examples as references helping instructors and teacher trainers make policy decisions according to their own teaching contexts. These templates typically showcase three distinct policy perspectives—restrictive, mixed, and encouraging, to accommodate diverse disciplines’ contexts. Harvard University’s syllabus samples [^3] https://oue.fas.harvard.edu/ai-guidance exemplify this approach (see Fig. 8). Universities usually encourage all instructors to explicitly include a clear policy in course syllabi regarding the use and misuse of ChatGPT and other GenAI tools. Open and explicit communications are crucial to help students understand the boundaries and expectations when they interact with GenAI tools in learning.
@@
Additionally, 41 (39.4%) universities include one-on-one consultations as a type of resource for instructors and/or students. They are conducted with the institution’s GenAI specialists and/or representatives from the Teaching and Learning Center to address attendees’ specific concerns and navigate the personal applications of ChatGPT and other GenAI tools in their own teaching and learning contexts.
Furthermore, it is evident in Fig. 7 that a smaller proportion of universities offer workshops (n = 40, 38.5%) and discussions (n = 20, 19.2%) regarding the use of ChatGPT and other GenAI tools. This might be because many of these workshops and discussions are a part of internal resources and are not publicly available. As ChatGPT is a new emerging technology, workshops and discussions are crucial for familiarizing educators with its positive applications and educational implications (De Winter et al., 2023). The results could reflect that there is a possible demand for increasing the frequency and accessibility of these events.
@@ 5.2.3 Content Analysis of Resources and Guidelines
Fig. 9 summarizes the focuses and purposes of the existing resources and guidelines regarding the use of ChatGPT and other GenAI tools in higher education. The majority of the universities start by introducing some general technical information (n = 85, 81.7%), such as a beginning section named What is ChatGPT? or GenAI in Education. This approach indicates the institutions’ intention to familiarize faculty, students, and staff with GenAI tools and enhance their understanding of these technologies. 62 (59.6%) universities discuss the ethical implications of implementing ChatGPT in higher education, including their apprehensions about the misuse of GenAI to foster plagiarism, violate academic integrity, and negatively affect student evaluation. This discussion highlights the necessity for instructors to monitor and guide the use of ChatGPT (Huallpa et al., 2023), as well as the need to adapt teaching methodologies to prevent possible cheating and plagiarism.
@@
56 (53.8%) universities explicitly list the inherent limitations of the current GenAI model and 38 (36.5%) universities address the data privacy issue, which serves as a reminder for individuals to keep in mind that these constraints may influence teaching approaches, learning experiences, and the conduction of research. Being aware of these limitations and issues ensures that GenAI tools would be used in a more responsible and ethical manner in education (Kasneci et al., 2023). The limitations primarily include:
@@
The inevitable intertwining with GenAI tools needs teachers and educators to evolve the teaching methods across all subjects. Therefore, numerous top 100 universities have offered resources on pedagogical applications (n = 56, 53.8%) and prevention strategies (n = 57, 54.8%). They cover the effective incorporation of ChatGPT into classrooms to enhance student learning experiences, alongside strategies to avoid its misuse by students. Table 4 provides a summary of major pedagogical applications and prevention strategies sourced from these leading universities. Both approaches highlight the significance of cultivating students’ critical thinking skills and problem-solving abilities, which are essential competencies in today’s rapidly advancing, technology-infused world.
@@
Pedagogical Applications Effectively incorporate AI in teaching and learning Prevention Strategies Prevent inappropriate use of AI in teaching and learning
Ask students to analyze and evaluate AI-generated texts. Ask students to compare and evaluate the different versions of texts generated by different AI tools. Ask students to compare/contrast AI-generated texts with human writing. Ask students to revise and edit AI-generated information. Ask students to debate or argue with AI and reflect on their learning. Use AI as a resource for students to receive feedback on their drafts Use AI tools to brainstorm initial teaching ideas and activities. Use AI tools to generate additional examples of certain concepts. Use AI tools to summarize long or difficult text. Use AI tools to generate writing prompts, grading rubrics or quiz questions based on the course materials. Ask students to explain their thought processes as they solve problems. Ask students to reflect on their personal learning experiences and opinions. Ask students to connect with their personal knowledge and life experiences. Ask students to include and provide proper academic citations. Ask students to reference class materials, notes, or sources that are unavailable online. Ask students to complete assignments in class. Ask students to present their answers in multimodal ways, such as hand drawing, or audio threads. Include visual prompts in assignments. Design assignments related to current events or discussions in the specific academic field. Divide the larger project into multiple smaller tasks.
@@
Another trend that emerges from the data is the discussion of using GenAI detection tools or GenAI detectors to identify AI-generated text in students’ work. 59 (56.7%) universities discuss the available common detective tools, such as Turnitin and GPTZero. However, it is worth noting that none (n = 0, 0%) of the universities in this study view the use of detective tools as a completely reliable method to identify AI-generated information, and none (n = 0, 0%) of them support instructors to use of the tools to evaluate students’ academic integrity and determine plagiarism. While GenAI detectors are designed to identify AI-generated language patterns, the research conducted by Sadasivan et al. (2023) shows that they are not reliable in many real-world scenarios. This is particularly evident when the detective tools are faced with paraphrasing attacks which refer to applying a light paraphraser to generated texts. Even a minor rephrasing can significantly affect the performance and accuracy of the entire detection system (Sadasivan et al., 2023). Additionally, universities in this study raise further concerns regarding the use of GenAI detectors, including the potential violation of students’ intellectual property rights and the risk to data privacy once their work is submitted to the detective tools. Some universities also believe that relying on such tools might undermine the relationship of trust between students and teachers as well.
5.3 Trends in Perceptions and Resource Provision of GenAI
RQ3 examines how universities’ perceptions and resources on the use of GenAI are affected by two possible factors: 1) institution ranking tiers; and 2) institution academic specializations. We will discuss the pedagogical implications that can be drawn from these findings.
5.3.1 Perception and Available Resources Across Different Tiers
This section employs the scale and point system described in method Section 4.3 to examine the relationship between three dimensions, including perceptual stances, resource diversity, and university rankings. We divide the top 100 universities (n = 104) into three tiers to examine the disparities in perceptions and resources across different tiers of higher education institutions. Tier 1 includes the top 1-33 universities (n = 34); Tier 2 comprises the subsequent top universities ranked 34-66 (n = 32); and Tier 3 includes the remaining reputational universities ranked 67-100 (n = 38).
In terms of the results, Fig. 10 illustrates the correlation between the perceptions and resources regarding GenAI across the three ranking tiers. We find that a majority of the top 100 universities have developed a diversity of resources and guidelines for integrating GenAI, demonstrating a proactive approach towards applying this technology to their education. However, the figure shows that there is no significant correlation between the universities’ rankings and the depth of perceptions and resources related to GenAI. The trend also indicates that a large number of universities present unclear and cautious perceptions toward GenAI. These perceptions might be due to uncertainties and controversial features of GenAI in higher education.